OpenStudy (anonymous):

Simplify: (7m^4)(2m^3)

OpenStudy (lgbasallote):

what do you think?

OpenStudy (jackellyn):

here's a hint @Danny17 when you multiply two variables with exponents the exponents add example: x(x^2)=x^3 because x=x^1 and 1+2=3. Can you apply this to your question?

OpenStudy (anonymous):

(7*2)m^7